40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 783     Entry time: Sat Aug 2 13:07:23 2008
Author: Koji 
Type: Configuration 
Category: General 
Subject: The AP table cleaned 
During the construction of the independent PLL I cleaned up some of the unused optics from the AP table. Essentially this should be harmless as they had already been isolated from any beam. They were related to Go's squeezing project and Osamu's MC Transmitted beam measurement.

Nevertherless, if you find any problem on the signals at the AP table (when the ifo returns), I am the person to be blamed.

I am going to update the table layout later next week.
ELOG V3.1.3-